OpenStudy (anonymous):

The coordinates of the vertices of quadrilateral ABCD are A(3,4), B(7,4), C(1,1), D(1,3). ABCD is reflected across the x-axis to form PQRS. What is the length of diagonal PR?

OpenStudy (anonymous):

Even though the figure is reflected, reflection is an isometry meaning that it doesn't change shape or size after the transformation. Therefore, PR's length is equivalent to AC because ABCD = PQRS. See how the letters are in the same position? Use the distance formula to find the distance. (3, 4) and (1, 1) \[d = \sqrt{(x_{1} - x_{2})^{2} + (y_{1} - y_{2})^{2}}\]\[d = \sqrt{(3 - 1)^{2} + (4 - 1)^{2}}\]Can you finish it up?

OpenStudy (anonymous):

\[\sqrt{13}\]

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

THANKU:)

OpenStudy (anonymous):

np ^_^